# A Hierarchical Framework for Measuring Scientific Paper Innovation via Large Language Models

### v1

@@ Abstract
Measuring scientific paper innovation is both important and challenging. Existing content-based methods often overlook the full-paper context, fail to capture the full scope of innovation, and lack generalization. We propose HSPIM, a hierarchical and training-free framework based on large language models (LLMs). It introduces a Paper-to-Sections-to-QAs decomposition to assess innovation. We segment the text by section titles and use zero-shot LLM prompting to implement section classification, question-answering (QA) augmentation, and weighted novelty scoring. The generated QA pair focuses on section-level innovation and serves as additional context to improve the LLM scoring. For each chunk, the LLM outputs a novelty score and a confidence score. We use confidence scores as weights to aggregate novelty scores into a paper-level innovation score. To further improve performance, we propose a two-layer question structure consisting of common and section-specific questions, and apply a genetic algorithm to optimize the question-prompt combinations. Comprehensive experiments on scientific conference paper datasets show that HSPIM outperforms baseline methods in effectiveness, generalization, and interpretability.
Index Terms:
Innovation Measurement, Hierarchical Framework, Text Mining, Large Language Model.
I Introduction
Recently, the rapid increase in scientific paper submissions has significantly heightened the demand for high-quality evaluation. The most important aspect of evaluating papers is to assess their innovation. Innovation was introduced in economics to explain how new products, processes, or structures emerge to transform industries and promote development [1]. In terms of research, innovation involves applying novel tools, methods, or concepts to enhance both methodologies and outcomes [2]. We note that innovation is different from novelty and originality. While novelty and originality focus on newness and uniqueness, they may lack sound results and practical contributions. Instead, innovation emphasizes both novelty and validated value [3, 4].
@@ I Introduction
Second, to measure innovation beyond novelty, we propose a weighted novelty scoring formula with question-answering (QA) augmentation. (i) We input each text chunk into a zero-shot LLM prompt to generate a JSON output {“novelty_score”, “reason”, “confidence_score”}. The novelty_score reflects section-level novelty, the reason explains the score, and the confidence_score indicates reliability. We then apply weighted averaging using the confidence scores to aggregate local novelty into a paper-level innovation score. (ii) We perform QA augmentation before scoring. Specifically, we design questions to guide the LLM to generate innovation-focused answers for each section chunk. The resulting QA pair is added as additional context to enhance in-context learning during LLM scoring. Inspired by structural and regulatory genes in genetics [11], we design a two-layer question structure: one common question shared across all sections, and one specific question for each section type. The generated QA pairs distill original section content into concise, innovation-focused carriers [12], helping the LLM analyze more aspects of innovation, such as feasibility and practical value. We refer to this Paper-to-Sections-to-QAs framework as Hierarchical Scientific Paper Innovation Measurement (HSPIM).
@@
Third, we use pre-trained LLMs as zero-shot generators and scorers to exploit their knowledge and reasoning strength. This approach is training-free and demonstrates strong generalization. To further improve robustness and effectiveness, we treat the question prompts as optimization targets in a combinatorial optimization task. Inspired by previous prompt optimization approaches [13, 14, 15] and gene variation theory, we propose the multi-prompt optimizer using the heuristic genetic algorithm (GA) [16]. We assume there are $s$ section types. GA searches for the optimal question-prompt combination within $s$ specific question sets and one common question set, following standard genetic algorithm steps: population initialization, selection, crossover, and mutation. Each individual in the population consists of $s$ specific questions plus one common question. Under the elite GA, we explore different optimization strategies for two-layer question structure, including joint optimization, two-step optimization, and pruning, to further enhance performance.
@@
We use three peer review datasets from PeerRead [17] and one dataset from NLPeer [18] to conduct extensive experiments. We average peer review scores for originality and soundness/correctness to form the ground-truth innovation score. Performance is evaluated using metrics such as RMSE and, around which we conduct further detailed analyses. Additionally, we evaluate the semantic textual similarity between reason output scoring reasons and actual peer review comments.
@@
  • We clearly define the concepts of innovation by distinguishing it from novelty and originality. Based on this distinction, we propose weighted novelty scoring and QA augmentation.
@@
II Related Work
II-A Innovation Concept
Schumpeter first introduced innovation in The Theory of Economic Development as creating new products, processes, or structures to disrupt equilibrium and drive growth [1]. Christensen later classified innovation into sustaining and disruptive types [19]. Disruptive innovation reshapes markets, gains value through adoption, and combines novelty with impact, making it hard to measure directly. In research, highly innovative work uses novel tools, methods, or concepts to improve both methodology and outcomes, arising from internal and external influences [2]. Note that innovation differs from novelty, which stresses uniqueness but may lack practical contribution [20, 21]. Prior studies indicate that innovation needs not only novelty but also contribution and value [3, 4, 22], so assessment must extend beyond novelty to capture real-world impact. A comprehensive analysis capturing the composite nature of innovation is essential before evaluation, as it cannot be directly measured like novelty.
II-B Approaches to Measuring Paper Innovation
@@
As previous studies suggest, paper novelty is evaluated by identifying atypical combinations of existing knowledge [23, 24]. Approaches to novelty measurement include content-based and metadata-based methods.
@@
Content-based methods rely on textual analysis to align with peer review practices that assess novelty directly from the text. These content-based models range from rule-based approaches to deep learning methods. Rule-based methods include WordNet [5] and domain-specific knowledge bases [25]. Deep learning models include unsupervised generative models such as graph-based autoencoder neural networks [6] and supervised discriminative models like convolutional neural network (CNN) [26], long short-term memory (LSTM) [27, 17], Transformer [28], TextCNN and BERTopic [29, 30]. Recently, fine-tuned LLM-based models like SEA [31] generate review texts along with several evaluation scores to simulate the real-world peer review. These content-based methods have a limited analysis scope, such as titles [29], abstracts [32], keywords [33], or contribution sentences [30]. This restriction may introduce biases [9] compared to full-text analysis.
@@
On the other hand, metadata-based methods evaluate external factors such as journal impact and author reputation through regression models [7]. Other metadata-based methods includes applying scientometrics techniques [8] and bibliometric methods [34, 35] based on reference analysis. Since content-based methods help to evaluate unpublished papers and support double-blind peer review, they are more practical than metadata-based methods for real-world applications. In addition, the innovation level of a paper does not depend on its external attributes.
@@
II-C Prompt-based LLMs
Large language models (LLMs) are effective tools for in-context learning [36, 37]. Zero-shot LLMs combine input data and prompt design to form context and then use this context to generate answers directly. Moreover, LLM-as-a-Judge [38, 39] shows that zero-shot LLMs can closely match human judgment and produce reliable scores. LLM-as-a-Judge can directly assign quantitative scores based on in-context learning. Thus, we suggest that zero-shot LLMs can effectively generate novelty and confidence scores for evaluating scientific paper innovation based on the paper content. Furthermore, we extract innovation-focused QA pairs from the original text chunk of the paper for better innovation judgment [12].
As previous studies have shown [40], LLM performance is sensitive to prompt designs. Therefore, prompt optimization has emerged as a prominent paradigm. Early works such as Optimization by PROmpting (OPRO) [14] and Automatic Prompt Optimization (APO) [41] leverage zero-shot LLMs as optimizers to update prompts based on output performance feedback. They apply heuristic algorithms to automate this process, enabling derivative-free and training-free optimization that surpasses manual designs. Other related methods include Evolutionary Prompt Optimizer (EvoPrompt) [13] and AutoHint [42].
In contrast to long-prompt optimization [15], which updates individual sentences within a large prompt in each iteration while maintaining the overall structure, our approach performs multi-prompt optimization. Instead of modifying a single long prompt, we need to jointly optimize multiple question prompts, each generating QA pairs for different text chunks. These QA pairs contribute to weighted novelty scoring to compute an innovation score, and we update question prompts based on RMSE performance feedback. To achieve this, we introduce a genetic algorithm-based method [16, 43] that simultaneously updates multiple prompts within HSPIM.
III Methodology
In this section, we first define the notations and the problem. Then, we provide a detailed introduction to the HSPIM framework and compare it with its simplified variant. Finally, we justify the two-layer question structure and propose three multi-prompt optimization strategies. A summary of key notations used in this paper is given in Table I.
III-A Notations and Problem Definition
@@
We derive ground-truth innovation scores from peer reviews in the scientific paper dataset. Each paper $p_{i}$ receives $m_{i}$ reviews $\{r_{i1},r_{i2},\dots,r_{im_{i}}\}$, each containing a comment and several evaluation scores. Following established peer review criteria, we quantify innovation using originality and soundness/correctness. Specifically, for each review $r_{ij}$, $\textit{Innovation}_{ij}=(\textit{Originality}_{ij}+\textit{Soundness}_{ij})/2$. The final labeled innovation score for $p_{i}$ is obtained by averaging its review scores: $\textit{Innovation}_{i}=\frac{1}{m_{i}}\sum_{j=1}^{m_{i}}\textit{Innovation}_{ij}$. Our objective is to minimize the discrepancy between the predicted innovation score $\hat{\textit{Innovation}}_{i}$ and the labeled innovation score $\textit{Innovation}_{i}$. To achieve this, we use root mean squared error (RMSE) defined as $\sqrt{\tfrac{1}{n}\sum_{i=1}^{n}(\hat{\textit{Innovation}}_{i}-\textit{Innovation}_{i})^{2}}$, where $n$ is the total number of papers.
TABLE I: Summary of notations in HSPIM framework
@@
$\textit{Confidence}_{ik}$ Predicted confidence score for chunk $t_{ik}$
$\textit{Novelty}_{ik}$ Predicted novelty score for chunk $t_{ik}$
@@
We propose the HSPIM framework to implement hierarchical section-based text mining. Each paper $p_{i}$ is divided into $l_{i}$ text chunks $\{t_{i1},t_{i2},\ldots,t_{il_{i}}\}$ and then classified into $s$ predefined section types $\{section_{1},section_{2},\dots,section_{s}\}$ by LLMs. By generating confidence scores and novelty scores for each section chunk via LLMs, the predicted innovation score $\hat{\textit{Innovation}}_{i}$ is obtained as a weighted average of the section novelty scores, with confidence scores serving as weights. Besides, HSPIM employs QA augmentation before scoring. It optimizes prompts to generate section-specific QA pairs. A genetic algorithm searches for the optimal prompt combination within $s$ predefined specific question sets $Q_{1},Q_{2},\dots,Q_{s}$ (one set per section type) and a single common question set $\mathcal{Q}_{c}$. Assume that $t_{ik}$ is classified as $section_{r}$ with $r\in\{1,2,\ldots,s\}$. For each text chunk $t_{ik}$, the final question $Q_{ik}$ is formed by concatenating one question from $Q_{r}$ and one question from $\mathcal{Q}_{c}$. Then $QA_{ik}$ is incorporated as additional innovation-focused information for $t_{ik}$ in the scoring prompt, guiding the LLM to generate more meaningful and informative scores. Here, we define our HSPIM as follows:
Definition III.1 (HSPIM).
@@
where $\textit{Confidence}_{ik}$ and $\textit{Novelty}_{ik}$ are LLM-generated values for $t_{ik}$. To enhance scoring, HSPIM applies QA augmentation and optimizes question prompts with genetic algorithm. It selects an optimal question-prompt combination from $s$ section-specific question sets and a shared common question set. For each chunk $t_{ik}$ classified as $section_{r}$, the final question $Q_{ik}$ is formed by combining one question from $Q_{r}$ and one from $\mathcal{Q}_{c}$. The resulting QA pairs provide additional context to guide the LLM in generating more reliable innovation scores.
III-B HSPIM Framework and Workflow

### v2

@@ Abstract
Measuring scientific paper innovation is both important and challenging. Existing content-based methods often overlook the full-paper context, fail to capture the full scope of innovation, and lack generalization. We propose HSPIM, a hierarchical and training-free framework based on large language models (LLMs). It introduces a Paper-to-Sections-to-QAs decomposition to assess innovation. We segment the text by section titles and use zero-shot LLM prompting to implement section classification, question-answering (QA) augmentation, and weighted innovation scoring. The generated QA pair focuses on section-level innovation and serves as additional context to improve the LLM scoring. For each chunk, the LLM outputs a novelty score and a confidence score. We use confidence scores as weights to aggregate novelty scores into a paper-level innovation score. To further improve performance, we propose a two-layer question structure consisting of common and section-specific questions, and apply a genetic algorithm to optimize the question-prompt combinations. Furthermore, under the fine-grained structure of innovation, we extend HSPIM to an HSPIM+ that generates novelty, contribution, and feasibility scores with respective confidence scores. Comprehensive experiments on scientific conference paper datasets show that HSPIM outperforms baseline methods in effectiveness, generalization, and interpretability. Demo code is available at https://github.com/Jasaxion/HSPIM.
keywords:
Innovation Measurement , Hierarchical Framework , Text Mining , Large Language Model
[…]
1 Introduction
Recently, the rapid increase in scientific paper submissions has significantly heightened the demand for high-quality evaluation. The most important aspect of evaluating papers is to assess their innovation. Innovation was introduced in economics to explain how new products, processes, or structures emerge to transform industries and promote development [1]. In terms of research, innovation involves applying novel tools, methods, or concepts to enhance both methodologies and outcomes [2]. We note that innovation is different from novelty. While novelty focuses on newness and uniqueness, it may lack practical feasibility and real-world contributions. Instead, innovation emphasizes both novelty and practical value [3].
@@ 1 Introduction
Second, to measure innovation beyond novelty, we follow social science theories to define it as a combination of novelty and practical value. Practical value includes contribution and feasibility. This structure reflects common understanding in innovation studies. Based on this innovation concept, we propose a weighted innovation scoring formula with question-answering (QA) augmentation. (i) We input each text chunk into a zero-shot LLM prompt to generate a JSON output {“novelty_score”, “reason”, “confidence_score”}. The novelty_score reflects section-level novelty. The reason explains the rationale behind the score. The confidence_score serves as a weight to modulate the novelty score of each section. Specifically, the confidence score reflects how strongly the LLM supports the local novelty claim in context. This design implicitly captures the feasibility attribute of innovation by indicating the plausibility of the novelty score. We then apply weighted averaging using the confidence scores to aggregate local novelty into a paper-level innovation score.
(ii) We perform QA augmentation before scoring. Specifically, we design questions to guide the LLM to generate innovation-focused answers for each section chunk. The resulting QA pair is added as additional context to enhance in-context learning during LLM scoring. Inspired by structural and regulatory genes in genetics, we design a two-layer question structure: one common question shared across all sections, and one specific question for each section type. The generated QA pairs distill original section content into concise, innovation-focused carriers [10], helping the LLM analyze more attributes of innovation, such as feasibility and contribution. We refer to this Paper-to-Sections-to-QAs framework as Hierarchical Scientific Paper Innovation Measurement (HSPIM). To evaluate the effect of a fine-grained representation of innovation, we further extend HSPIM to HSPIM+ by incorporating a $p$-norm aggregation of novelty, contribution, and feasibility scores, each with corresponding confidence weights.
@@
Third, we use pre-trained LLMs as zero-shot generators and scorers to exploit their knowledge and reasoning strength. This approach is training-free and demonstrates strong generalization. To further improve robustness and effectiveness, we treat the question prompts as optimization targets in a combinatorial optimization task. Inspired by previous prompt optimization approaches [11, 12] and gene variation theory, we propose the multi-prompt optimizer using the heuristic genetic algorithm (GA) [13]. We assume there are $s$ section types. GA searches for the optimal question-prompt combination within $s$ specific question sets and one common question set, following standard genetic algorithm steps: population initialization, selection, crossover, and mutation. Each individual in the population consists of $s$ specific questions plus one common question. Under the elite GA, we explore different optimization strategies for two-layer question structure, including joint optimization, two-step optimization, and pruning, to further enhance performance.
@@
We use three peer review datasets from PeerRead [14] and one dataset from NLPeer [15] to conduct extensive experiments. We average peer review scores for originality and soundness/correctness as proxies for novelty and practical value, respectively, to construct the ground-truth innovation score. Performance is evaluated using metrics such as RMSE and MAE, around which we conduct further detailed analyses. Additionally, we evaluate the semantic textual similarity between reason output scoring reasons and actual peer review comments.
In addition to traditional content-based approaches, recent LLM-based models such as SEA-E [16] focus on generating peer reviews with attributes like originality and soundness. However, these systems are typically fine-tuned on specific datasets, operate in a single-pass manner, and capture review-style criteria rather than directly measuring innovation. In contrast, our framework represents the first LLM-based approach for scientific innovation measurement, realized through a training-free, hierarchical evaluation that integrates section-level QAs with a genetic multi-prompt optimizer.
@@
  • 2.
    We clearly define the concepts of innovation by distinguishing it from novelty. Based on this distinction, we propose weighted innovation scoring and QA augmentation.
  • 3.
@@
2 Related Work
Our work draws on prior research on three areas: the concept of innovation, paper innovation measurement, and prompt-based LLMs.
2.1 Innovation Concept
Schumpeter initially defined innovation in The Theory of Economic Development as creating new products, processes, or structures to disrupt equilibrium and drive growth [1]. Christensen further distinguished innovation into sustaining and disruptive types, emphasizing that disruptive innovation combines novelty with real-world impact through market reshaping and widespread adoption [17]. These classic economic theories highlight that innovation inherently focuses on creating value through real-world application and impact.
Social science research provides multifaceted insights into the nature of innovation. Early work highlights that innovation should respond to real-world needs through public demand and institutional frameworks. Later studies suggest that innovation goes beyond novelty, encompassing practical value within broader ecological and methodological contexts [18]. Other research shows that innovation is socially constructed and requires recognition from audiences and institutions. Recent work in research assessment suggests that innovation involves both methodological novelty and impact-oriented contributions [2]. Together, these studies support the view that innovation entails both novelty and practical value.
Historically, innovation evaluation focused mainly on novelty. Over time, it began to include practical value, such as applicability and feasibility, recognizing that innovation is a multifaceted concept. Dosi introduced the concepts of technological paradigms and technological trajectories, clarifying that innovation involves not just novelty, but also systematic realizability and practical feasibility [19]. Similarly, Rogers et al. argued that innovation diffusion depends on two key factors. The first is relative advantage, which reflects real-world contribution. The second is complexity and trialability, which reflect feasibility [20].
Figure 1: Conceptual decomposition of innovation into novelty and practical value (contribution and feasibility), which grounds the scoring design of HSPIM.
Integrating these insights, we define innovation as a composite concept with two main components: Novelty and Practical Value. Practical value is further divided into Contribution (theoretical, methodological, or applied value) and Feasibility (the potential for real-world implementation), as shown in Fig. 1. Since innovation is a complex concept, it cannot be directly measured like novelty. Therefore, a comprehensive analysis is needed to capture its full structure before quantification.
2.2 Approaches to Measuring Paper Innovation
@@
Existing methods for measuring paper innovation mainly focus on novelty. As previous studies suggest, paper novelty is evaluated by identifying atypical combinations of existing knowledge [21, 22]. Approaches to novelty measurement include content-based and metadata-based methods.
@@
Content-based methods rely on textual analysis to align with peer review practices that assess novelty directly from the text. These content-based models range from rule-based approaches to deep learning methods. Rule-based methods include WordNet [4] and domain-specific knowledge bases [23]. Deep learning models include unsupervised generative models such as graph-based autoencoder neural networks [5] and supervised discriminative models like convolutional neural network (CNN) [24], long short-term memory (LSTM) [25, 14], Transformer [26], TextCNN and BERTopic [27, 28]. Besides, recent studies have been conducted on text novelty detection tasks [29]. Recently, fine-tuned LLM-based models like SEA [16] generate review texts along with several evaluation scores to simulate the real-world peer review.
Content-based methods can be categorized by their analysis target, such as topics [30], titles [27], keywords [31], paragraphs, or contribution sentences [28]. Focusing on specific parts of the paper may introduce biases [8] compared to using full-text analysis.
@@
On the other hand, metadata-based methods evaluate external factors such as journal impact and author reputation through regression models [6]. Other metadata-based methods includes applying scientometrics techniques [7] and bibliometric methods [32] based on reference analysis. Since content-based methods help to evaluate unpublished papers and support double-blind peer review, they are more practical than metadata-based methods for real-world applications. In addition, the innovation level of a paper does not depend on its external attributes.
@@
2.3 Prompt-based LLMs
Large language models (LLMs) are effective tools for in-context learning [33]. Zero-shot LLMs combine input data and prompt design to form context and then use this context to generate answers directly. Moreover, LLM-as-a-Judge [34] shows that zero-shot LLMs can closely match human judgment and produce reliable scores. LLM-as-a-Judge can directly assign quantitative scores based on in-context learning. Thus, we suggest that zero-shot LLMs can effectively generate novelty and confidence scores for evaluating scientific paper innovation based on the paper content. Furthermore, we extract innovation-focused QA pairs from the original text chunk of the paper for better innovation judgment [10].
As previous studies have shown [35], LLM performance is sensitive to prompt designs. Therefore, prompt optimization has emerged as a prominent paradigm. Early works such as Optimization by PROmpting (OPRO) [12] and Automatic Prompt Optimization (APO) leverage zero-shot LLMs as optimizers to update prompts based on output performance feedback. They apply heuristic algorithms to automate this process, enabling derivative-free and training-free optimization that surpasses manual designs. Other related methods include Evolutionary Prompt Optimizer (EvoPrompt) [11] and AutoHint [36].
In contrast to long-prompt optimization, which updates individual sentences within a large prompt in each iteration while maintaining the overall structure, our approach performs multi-prompt optimization. Instead of modifying a single long prompt, we need to jointly optimize multiple question prompts, each generating QA pairs for different text chunks. These QA pairs contribute to weighted innovation scoring to compute an innovation score, and we update question prompts based on RMSE performance feedback. To achieve this, we introduce a genetic algorithm-based method [13] that simultaneously updates multiple prompts within HSPIM.
3 Methodology
In this section, we first define the notations and the problem. Then, we provide a detailed introduction to the HSPIM framework and compare it with its simplified variant. Finally, we justify the two-layer question structure and propose three multi-prompt optimization strategies. A summary of key notations used in this paper is given in Table 1.
3.1 Notations and Problem Definition
@@
We derive ground-truth innovation scores from peer reviews in the scientific paper dataset. Each paper $p_{i}$ receives $m_{i}$ reviews $\{r_{i1},r_{i2},\dots,r_{im_{i}}\}$, each containing a comment and several evaluation scores. Following established peer review criteria, we quantify innovation using originality and soundness/correctness. Specifically, for each review $r_{ij}$, the innovation score is defined as $\textit{Innovation}_{ij}=(\textit{Originality}_{ij}+\textit{Soundness}_{ij})/2$. The final labeled innovation score for $p_{i}$ is obtained by averaging its review scores: $\textit{Innovation}_{i}=\frac{1}{m_{i}}\sum_{j=1}^{m_{i}}\textit{Innovation}_{ij}$. Our objective is to minimize the discrepancy between the predicted innovation score $\hat{\textit{Innovation}}_{i}$ and the labeled innovation score $\textit{Innovation}_{i}$. To achieve this, we use the root mean squared error (RMSE), calculated as the square root of the average of $(\hat{\textit{Innovation}}_{i}-\textit{Innovation}_{i})^{2}$ over all $n$ papers.
Table 1: Summary of notations in HSPIM framework.
@@
$\textit{Confidence}_{ik}$ ($c_{ik}$) Predicted confidence score for chunk $t_{ik}$
$\textit{Novelty}_{ik}$ ($n_{ik}$) Predicted novelty score for chunk $t_{ik}$
$\textit{Contribution}_{ik}$ ($d_{ik}$) Predicted contribution score for chunk $t_{ik}$
$\textit{Feasibility}_{ik}$ ($f_{ik}$) Predicted feasibility score for chunk $t_{ik}$
@@
We propose the HSPIM framework to implement hierarchical section-based text mining. Each paper $p_{i}$ is divided into $l_{i}$ chunks $\{t_{i1},t_{i2},\ldots,t_{il_{i}}\}$, which are classified into $s$ predefined section types $\{section_{1},\dots,section_{s}\}$ by LLMs. By generating confidence scores and novelty scores for each section chunk via LLMs, the predicted innovation score $\hat{\textit{Innovation}}_{i}$ is obtained as a weighted average of the section novelty scores, with confidence scores serving as weights. Besides, HSPIM employs QA augmentation before scoring. It optimizes prompts to generate section-specific QA pairs. A genetic algorithm searches for the optimal prompt combination within $s$ predefined specific question sets $Q_{1},Q_{2},\dots,Q_{s}$ (one set per section type) and a single common question set $\mathcal{Q}_{c}$. Assume that $t_{ik}$ is classified as $section_{r}$ with $r\in\{1,2,\ldots,s\}$. For each text chunk $t_{ik}$, the final question $Q_{ik}$ is formed by concatenating one question from $Q_{r}$ and one question from $\mathcal{Q}_{c}$. Then $QA_{ik}$ is incorporated as additional innovation-focused information for $t_{ik}$ in the scoring prompt, guiding the LLM to generate more meaningful and informative scores. Here, we define our HSPIM as follows:
Definition 3.1 (HSPIM).
@@
where $\textit{Confidence}_{ik}$ and $\textit{Novelty}_{ik}$ are LLM-generated values for $t_{ik}$. To enhance scoring, HSPIM uses QA augmentation and a genetic algorithm to optimize question prompts. It selects an optimal question-prompt combination from $s$ section-specific question sets and a shared common question set. For each chunk $t_{ik}$ classified as $section_{r}$, the final question $Q_{ik}$ is formed by combining one question from $Q_{r}$ and one from $\mathcal{Q}_{c}$. The resulting QA pairs provide additional context to guide the LLM in generating more reliable innovation scores.
We also introduce HSPIM+ as a norm-based extension of HSPIM. The $p$-norm allows us to combine novelty, contribution, and feasibility in a flexible way. Different values of $p$ let the final score reflect a balanced, cumulative, or dominant effect.
Definition 3.2 (HSPIM+ (Norm-based Extension)).
HSPIM+ extends the HSPIM framework in Definition 3.1 by modeling innovation as a composite of three attributes: novelty, contribution, and feasibility. For each chunk $t_{ik}$, LLMs generate $\textit{Novelty}_{ik}$, $\textit{Contribution}_{ik}$, and $\textit{Feasibility}_{ik}$, along with corresponding confidence scores $c^{(n)}_{ik}$, $c^{(d)}_{ik}$, $c^{(f)}_{ik}$ as weights.
First, for each attribute, a weighted mean is computed:
$\textit{Novelty}_{i}$ $=\frac{\sum_{k=1}^{l_{i}}c^{(n)}_{ik}\cdot\textit{Novelty}_{ik}}{\sum_{k=1}^{l_{i}}c^{(n)}_{ik}}$ (2)
$\textit{Contribution}_{i}$ $=\frac{\sum_{k=1}^{l_{i}}c^{(d)}_{ik}\cdot\textit{Contribution}_{ik}}{\sum_{k=1}^{l_{i}}c^{(d)}_{ik}}$ (3)
$\textit{Feasibility}_{i}$ $=\frac{\sum_{k=1}^{l_{i}}c^{(f)}_{ik}\cdot\textit{Feasibility}_{ik}}{\sum_{k=1}^{l_{i}}c^{(f)}_{ik}}$ (4)
Let $\mathbf{v}_{i}=[\textit{Novelty}_{i},\,\textit{Contribution}_{i},\,\textit{Feasibility}_{i}]$ be the score vector for paper $p_{i}$. The overall innovation score is then given by the $p$-norm of this vector, linearly mapped back to the $[1,5]$ range:
$\hat{\textit{Innovation}}_{i}^{(p)}=\mathcal{M}\left(\left\|\mathbf{v}_{i}\right\|_{p}\right)$ (5)
where $\mathcal{M}(\cdot)$ denotes a linear mapping to $[1,5]$, and $p\in\{1,2,\infty\}$ specifies the aggregation norm: $L_{1}$ (mean), $L_{2}$ (Euclidean), and $L_{\infty}$ (max), respectively. Here, the confidence scores reflect the credibility of the LLMs’ evaluation of novelty, contribution, and feasibility.
3.2 HSPIM Framework and Workflow